# Certifying ground-state properties of quantum many-body systems

### v4

@@ II Ground-state problem
We consider quantum systems composed of $N$ particles, whose interactions are described by a Hamiltonian operator. In what follows, and for simplicity, we are going to focus our discussion on systems of finite dimension $d$ on a lattice, although the techniques we discuss also apply beyond this scenario, e.g. to boson or fermion models. We consider systems with local interactions so that it is possible to define a Hamiltonian for an arbitrary number of parties $H_{N}$ given by a linear combination of different tensor product terms $h_{i}$ acting on a subset of neighbouring particles, $H_{N}=\sum_{i}a_{i}h_{i}$. One is then interested in determining the value $o_{\text{GS}}^{(N)}$ of relevant physical observables $O_{N}$ in the ground state, $|\psi_{\text{GS}}^{(N)}\rangle$, see also Eq. 1, namely $o_{\text{GS}}^{(N)}=\langle\psi_{\text{GS}}^{(N)}|O_{N}|\psi_{\text{GS}}^{(N)}\rangle$. Often, one is also interested in the thermodynamic limit of an infinite number of particles, which are typically inferred using scaling considerations by studying the dependence of $o_{\text{GS}}^{(N)}$ with $N$. In what follows, we often remove the dependence on $N$ to simplify the notation.
@@ IV.1 Algorithmic considerations
Again, this is because the Hamiltonian can be expressed as a polynomial over the Pauli matrices at each site, $\{\sigma_{i}^{a}\}_{i=1,\ldots,N;a\in\{x,y,z\}}$, that is, as a liner combination of some monomials over operators of the form $u_{m}=\sigma_{i_{1}}^{a_{1}}\dots\sigma_{i_{p_{m}}}^{a_{p_{m}}}$, where $p_{m}$ is the degree of the monomial, having $H(\{\sigma_{i}^{a}\})=\sum_{m}c_{m}u_{m}$. As explained above, the problem can equivalently be seen as a minimisation of the linear function $\langle H\rangle=\sum_{m}c_{m}\langle u_{m}\rangle$ over the set of quantum (Pauli) moments ${\mathcal{M}}_{Q}=\{\{\langle u_{i}\rangle\}_{i}:\exists|\psi\rangle\text{ such that }\forall i~{}\langle u_{i}\rangle=\langle\psi|u_{i}|\psi\rangle\}$.
Without entering into the details, the SDP relaxations of [8, 9] replace the set of quantum moments ${\mathcal{M}}_{Q}$ in the optimisation by larger moment sets ${\mathcal{M}}$, hence providing lower bounds of the ground-state energy in Eq. 8. This is achieved by introducing the so-called moment matrices. Specifically, suppose that =${\mathcal{B}}=\{u_{m}\}_{m}$ is a (Pauli) monomial list. For each quantum realisation (namely: for any given quantum state), the quantum moment matrix $\mathbf{M}$ indexed by ${\mathcal{B}}$ is defined as $[{\mathbf{M}}]_{uv}=\langle u^{\dagger}v\rangle$. The entries of these quantum moment matrices satisfy a series of linear relations resulting from the constraints in the optimisation, in our case coming from the Pauli algebra and the commutation of operators acting at different sites. It is rather easy to see that, for each choice of the monomial list ${\mathcal{B}}$, the corresponding quantum moment matrix is positive, namely, $\mathbf{M}\succeq 0$. However, the opposite is not necessarily true: there exist positive moment matrices satisfying the linear relations associated to the constraints that do not have a quantum realisation. The different lists of monomials, and corresponding moment matrices, define the different sets of moments ${\mathcal{M}}$ over which the relaxations are built. For that, it suffices to consider monomial lists such that the energy can be expressed as a linear combination of the entries of the corresponding moment matrix:
@@
$\min\limits_{\{\langle u^{\dagger}v\rangle\}_{u,v\in{\mathcal{B}}}}$ $\quad\hbox{\rm{tr}}(H_{M}\mathbf{M})$ (10)
@@ IV.2 Heisenberg chain
Ground-state energy.– The ground-state energy per spin is given by $e_{\rm PBC}(N)=\langle H\rangle/N=3C_{1}$. In Fig. 2, we plot the best lower bound of $e_{\rm PBC}$ as obtained by our SDP relaxation for up to $N=100$ spins. As a comparison, we plot the (quasi-)exact energy as obtained by DMRG simulations. It is also of interest to compare our SDP lower bound with the Anderson bound, which is obtained by exact diagonalisation on a system with open boundary conditions (OBC), $e_{\rm OBC}(K)\leq e_{\rm PBC}(N)$ for all $N>K$ (see Appendix A for details on the Anderson bound). In Fig. 2, the SDP bound is seen to vastly outperform the Anderson bound (which is in fact estimated with DMRG for the sake of illustration for up to $N=100$, as beyond a few tens of qubits exact diagonaisation is out of reach).
@@ IV.3 Heisenberg chain with second-neighbour couplings
We then study the ability of the SDP approach to bound the spin correlation function at larger distance. In particular, as mentioned, for values of $J_{2}<J_{2,c}$, one expects that the system develops antiferromagnetic (that is, staggered) spin correlations which decay as a power-law with distance, as in the $J_{2}=0$ limit [42]. In order to explore the potentiality of SDP relaxations to capture such quasi-long-range order the ground state, we compute bounds on the spin-spin correlations as a function of distance for a fixed system size of $N=40$. We consider both $J_{2}=0.2<J_{2,c}$ (Fig. 6 and Table 7) and $J_{2}=1.0>J_{2,c}$ (Fig. 7 and Table 8). As can be seen, the SDP upper and lower bounds tightly sandwich the DMRG value at small distances, while they become looser at larger distances. Yet, one sees that
@@ IV.4 Square lattice Heisenberg model
where $(i,j)$ label the position of the spins on a square lattice with PBC ($L+1\equiv 1$). In contrast to the 1D model, it is expected that the square-lattice Heisenberg model spontaneously breaks the $SU(2)$ symmetry in the thermodynamic limit and displays true long-range antiferromagnetic order in the ground state, implying in particular that $C(L/2,L/2)\to{\rm cst.}>0$ for $L\to\infty$. As further discussed below, while we do certify this property for $L\leq 8$, we cannot reliably extrapolate the obtained SDP bounds to the thermodynamic limit.
@@ V Venues for improvements
Optimality constraints: The solution to optimisation problems have additional constraints that can be incorporated to the hierarchy. For instance, in many situations, such as for energy minimisation, the searched state is an eigenstate of the polynomial to be optimised, the Hamiltonian in our case. This can be used to enforce additional polynomial constraints that are satisifed by the ground state, e.g. $\langle[H,O]\rangle_{\text{GS}}=0$ for any operator $O$. In this sense, recent works, [57] and [58], appearing after the completion of a first version of the present manuscript, provide other ground-state relaxations that exploit the optimality constraints. Interestingly, these constraints allow deriving certifiable bounds on the value of any observable without the need for upper bounds on the energy. Ref. [58] is in fact more general and shows how to extend the Karush-Kuhn-Tucker optimality conditions, well known in classical optimisation, to the non-commuting case. It is also reported in [59] that adding entropy constraints could strengthen the SDP bounds on ground-state problems. Note that many of these constraints are in principle not needed because recovered from the convergence of the hierarchy. However, in practical terms, they may significantly improve the results at finite steps of the relaxation.
@@ VII Acknowledgements
We thank Miguel Navascués, Luca Tagliacozzo and Filippo Vicentini for discussions. This work is supported by the ERC AdG CERQUTE, the Government of Spain (NextGenerationEU PRTR-C17.I1 and Quantum in Spain, Severo Ochoa CEX2019-000910-S), Fundació Cellex, Fundació Mir-Puig, Generalitat de Catalunya (CERCA programme), the AXA Chair in Quantum Information Science, EU Quantera project Veriqtas and PASQUANS2, the National Natural Science Foundation of China under Grant No. 12201618, the European Union Horizon’s 2020 research and innovation programme under the Marie Sklodowska Curie grant agreement No 101031549 (QuoMoDys), the NSF grant OAC-1835443 and the ERC Adv. Grant grant agreement No 885682. Research at the Perimeter Institute for Theoretical Physics is supported by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Research, Innovation and Science. VM was supported by the FastQI grant funded by the Institut Quantique Occitan, the PHC Proteus grant 46195TA, the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Actions, grant agreement 813211 (POEMA), by the AI Interdisciplinary Institute ANITI funding, through the French “Investing for the Future PIA3” program under the Grant agreement ${}^{\circ}$ ANR-19-PI3A-0004 as well as by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. This work was partially performed using HPC resources from CALMIP (Grant 2023-P23035).
@@ Appendix B SDP reductions by exploiting structure
In this appendix, we further discuss how to implement symmetries of the problem in the SDP algorithm to reduce the number of free variables in the implementation. We start be restating the main problem for completeness.
@@
such that: $\quad(\sigma_{i}^{a})^{2}=1,\quad i=1,\ldots,N,a\in\{x,y,z\},$
@@
$\quad\sigma_{i}^{a}\sigma_{j}^{b}=\sigma_{j}^{b}\sigma_{i}^{a},\quad 1\leq i\neq j\leq N,a,b\in\{x,y,z\}.$
@@
$\min\limits_{\{\langle u^{\dagger}v\rangle\}}$ $\quad\langle H\rangle$ (18)
@@
where $\mathbf{M}_{d}$ is the moment matrix indexed by ${\mathcal{B}}_{d}$ with $[{\mathbf{M}}_{d}]_{uv}=\langle u^{\dagger}v\rangle$. The decision variables of (18) are the moments $\langle u^{\dagger}v\rangle,u,v\in{\mathcal{B}}_{d}$. Note that the equality constraints in (17) give rise to the following replacement rules on monomials:
@@ B.2.2 Sign symmetry of the Hamiltonian
Besides the zero entries given in Section B.2.1, these additional sign symmetries of the Hamiltonian yield extra zero entries of the moment matrix: $\langle u\rangle=0$ if $\hbox{\rm{NF}}(u)$ is variant under the transformations (B.2.2).
Note that symmetry reduction usually applies to any convex problem with symmetric objective and symmetric feasible set. In this case, the symmetry of the feasible set is not apparent because the replacement rules are not symmetric, e.g., replacing $\sigma_{i}^{x}\sigma_{i}^{y}$ by $\hbox{\bf{i}}\sigma_{i}^{z}$ is not symmetric under any of the substitutions (B.2.2). We can show the symmetry of the set of sums of Hermitian squares using (1) their connection with the set of strictly positive Hermitian elements over the quotient ring (detailed below) and (2) the fact that the Hamiltonian is invariant under the substitutions (B.2.2).
Indeed, let us denote the non-commutative polynomial ring by ${\mathbb{C}}\langle\{\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z}\}_{i=1}^{N}\rangle$ and the ideal generated by the equality constraints of (17) by $I$. Let $\Omega\coloneqq\{\mathrm{NF}(u)\mid u\text{ is a monomial in }\{\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z}\}_{i=1}^{N}\}$. Then the optimization problem (17) is equivalent to the unconstrained optimization problem: $\min\limits_{\{|\psi\rangle,\sigma_{i}^{a}\}}\langle\psi|H|\psi\rangle$, considered in the quotient ring ${\mathbb{C}}\langle\{\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z}\}_{i=1}^{N}\rangle/I\cong{\mathbb{C}}\langle\Omega\rangle$. Let us denote by $S$ the group of additional sign symmetries given by (B.2.2) and by $\Sigma_{S}$ the set of sums of Hermitian squares of ${\mathbb{C}}\langle\Omega\rangle$, that are invariant under $S$ after conversion to normal form, i.e., elements of the form $h=\sum_{j}p_{j}^{\dagger}p_{j}$, $p_{j}\in{\mathbb{C}}\langle\Omega\rangle$, such that $s(\hbox{\rm{NF}}(h))=\hbox{\rm{NF}}(h)$ for any $s\in S$. Then one can show that any strictly positive Hermitian element of ${\mathbb{C}}\langle\Omega\rangle$ that is invariant under $S$ lies in $\Sigma_{S}$ by [10], the proof being very similar to the one of [71, Proposition 3.1]. Here “strict positivity” should be understood as strict positivity over all possible evaluations in Hilbert spaces, as detailed, e.g., in [71, Section 2.2]. By duality one considers optimization over linear functionals nonnegative on $\Sigma_{S}$, which leads to an equivalent formulation of the above unconstrained optimization problem:
@@
To conclude, at each relaxation (18) we can restrict ourselves to optimizing over linear functionals vanishing on variant elements of ${\mathbb{C}}\langle\Omega\rangle$ under the transformations (B.2.2). This boils down to setting every entry of the Hermitian moment matrix ${\mathbf{M}}=[\ell(v^{\dagger}w)]_{v,w\in\Omega}$ from (18) to $\langle u\rangle=0$ if $\hbox{\rm{NF}}(u)$ is variant under the transformations (B.2.2).

### v5

@@ II Ground-state problem
We consider quantum systems composed of $N$ particles, whose interactions are described by a Hamiltonian operator. In what follows, and for simplicity, we are going to focus our discussion on spin systems of finite dimension $d$ on a lattice. The techniques we discuss also apply beyond this scenario, e.g. to fermions or boson models, once one takes into account the respective commutation-type properties of the creation and annihilation operators. We consider systems with local interactions so that it is possible to define a Hamiltonian for an arbitrary number of particles $H_{N}$ given by a linear combination of different tensor product terms $h_{i}$ acting on a subset of neighbouring particles, $H_{N}=\sum_{i}a_{i}h_{i}$. One is then interested in determining the value $o_{\text{GS}}^{(N)}$ of relevant physical observables $O_{N}$ in the ground state, $|\psi_{\text{GS}}^{(N)}\rangle$, see also Eq. 1, namely $o_{\text{GS}}^{(N)}=\langle\psi_{\text{GS}}^{(N)}|O_{N}|\psi_{\text{GS}}^{(N)}\rangle$. Often, one is also interested in the thermodynamic limit of an infinite number of particles, which are typically inferred using scaling considerations by studying the dependence of $o_{\text{GS}}^{(N)}$ with $N$. In what follows, we often remove the dependence on $N$ to simplify the notation.
@@ IV.1 Algorithmic considerations
Again, this is because the Hamiltonian can be expressed as a polynomial over the Pauli matrices at each site, $\{\sigma_{i}^{a}\}_{i=1,\ldots,N;a\in\{x,y,z\}}$, that is, as a linear combination of some monomials over operators of the form $v_{m}=\sigma_{i_{1}}^{a_{1}}\dots\sigma_{i_{p_{m}}}^{a_{p_{m}}}$, where $p_{m}$ is the degree of the monomial, having $H(\{\sigma_{i}^{a}\})=\sum_{m}c_{m}v_{m}$. As explained above, the problem can equivalently be seen as a minimisation of the linear function $\langle H\rangle=\sum_{m}c_{m}\langle v_{m}\rangle$ over the set of quantum (Pauli) moments ${\mathcal{M}}_{Q}=\{\{\langle v_{i}\rangle\}_{i}:\exists|\psi\rangle\text{ such that }\forall i~{}\langle v_{i}\rangle=\langle\psi|v_{i}|\psi\rangle\}$.
Without entering into the details, the SDP relaxations of [8, 9] replace the set of quantum moments ${\mathcal{M}}_{Q}$ in the optimisation by larger moment sets ${\mathcal{M}}$, hence providing lower bounds of the ground-state energy in Eq. 8. This is achieved by introducing the so-called moment matrices. Specifically, suppose that =${\mathcal{B}}=\{v_{m}\}_{m}$ is a (Pauli) monomial list. For each quantum realisation (namely: for any given quantum state), the quantum moment matrix $\mathbf{M}$ indexed by ${\mathcal{B}}$ is defined as $[{\mathbf{M}}]_{vw}=\langle v^{\dagger}w\rangle$. The entries of these quantum moment matrices satisfy a series of linear relations resulting from the constraints in the optimisation, in our case coming from the Pauli algebra and the commutation of operators acting at different sites. It is rather easy to see that, for each choice of the monomial list ${\mathcal{B}}$, the corresponding quantum moment matrix is positive, namely, $\mathbf{M}\succeq 0$. However, the opposite is not necessarily true: there exist positive moment matrices satisfying the linear relations associated to the constraints that do not have a quantum realisation. The different lists of monomials, and corresponding moment matrices, define the different sets of moments ${\mathcal{M}}$ over which the relaxations are built. For that, it suffices to consider monomial lists such that the energy can be expressed as a linear combination of the entries of the corresponding moment matrix:
@@
$\min\limits_{\{\langle v^{\dagger}w\rangle\}_{v,w\in{\mathcal{B}}}}$ $\quad\hbox{\rm{tr}}(H_{M}\mathbf{M})$ (10)
@@ IV.2 Heisenberg chain
Ground-state energy.– The ground-state energy per spin is given by $e_{\rm PBC}(N)=\langle H\rangle/N=3C_{1}$. In Fig. 2, we plot the best lower bound of $e_{\rm PBC}$ as obtained by our SDP relaxation for up to $N=100$ spins. As a comparison, we plot the (quasi-)exact energy as obtained by DMRG simulations. It is also of interest to compare our SDP lower bound with the Anderson bound, which is obtained by exact diagonalisation on a system with open boundary conditions (OBC), $e_{\rm OBC}(K)\leq e_{\rm PBC}(N)$ for all $N>K$ (see Appendix A for details on the Anderson bound). In Fig. 2, the SDP bound is seen to vastly outperform the Anderson bound (which is in fact estimated with DMRG for the sake of illustration for up to $N=100$, as beyond a few tens of qubits exact diagonalisation is out of reach).
@@ IV.3 Heisenberg chain with second-neighbour couplings
We then study the ability of the SDP approach to bound the spin correlation function at larger distance. In particular, as mentioned, for values of $J_{2}<J_{2,c}$, one expects that the system develops antiferromagnetic (that is, staggered) spin correlations which decay as a power-law with distance, as in the $J_{2}=0$ limit [42]. In order to explore the potentiality of SDP relaxations to capture such quasi-long-range order in the ground state, we compute bounds on the spin-spin correlations as a function of distance for a fixed system size of $N=40$. We consider both $J_{2}=0.2<J_{2,c}$ (Fig. 6 and Table 7) and $J_{2}=1.0>J_{2,c}$ (Fig. 7 and Table 8). As can be seen, the SDP upper and lower bounds tightly sandwich the DMRG value at small distances, while they become looser at larger distances. Yet, one sees that
@@ IV.4 Square lattice Heisenberg model
where $(i,j)$ label the position of the spins on a square lattice with PBC ($L+1\equiv 1$). In contrast to the 1D model, it is expected that the square-lattice Heisenberg model spontaneously breaks the $SU(2)$ symmetry in the thermodynamic limit and displays true long-range antiferromagnetic order in the ground state. In particular, this implies that $C(L/2,L/2)\to{\rm cst.}>0$ for $L\to\infty$, where we define $C(i,j)=(1/4)\langle\sigma_{(0,0)}^{x}\sigma_{(i,j)}^{x}\rangle$. As further discussed below, while we do certify this property for $L\leq 8$, we cannot reliably extrapolate the obtained SDP bounds to the thermodynamic limit.
@@ V Venues for improvements
Optimality constraints: The solution to optimisation problems have additional constraints that can be incorporated to the hierarchy. For instance, in many situations, such as for energy minimisation, the searched state is an eigenstate of the polynomial to be optimised, the Hamiltonian in our case. This can be used to enforce additional polynomial constraints that are satisifed by the ground state, e.g. $\langle[H,O]\rangle_{\text{GS}}=0$ for any operator $O$. In this sense, recent works, [57] and [58], appearing after the completion of a first version of the present manuscript, provide other ground-state relaxations that exploit the optimality constraints. Interestingly, these constraints allow deriving certifiable bounds on the value of any observable without the need for upper bounds on the energy. Ref. [58] is in fact more general and shows how to extend the Karush-Kuhn-Tucker optimality conditions, well known in classical optimisation, to the non-commuting case. It is also reported in [59] that adding entropy constraints could strengthen the SDP bounds on ground-state problems. Note that as the SDP hieararchy, when increasing the relaxation order, converges to the exact ground-state solution where these constraints are automatically satisfied, such additional constraints are in principle not needed. However, in practice, imposing them may significantly improve the results at finite steps of the relaxation.
@@ VII Acknowledgements
We thank Miguel Navascués, Luca Tagliacozzo and Filippo Vicentini for discussions. This work is supported by the ERC AdG CERQUTE, the Government of Spain (NextGenerationEU PRTR-C17.I1 and Quantum in Spain, Severo Ochoa CEX2019-000910-S), Fundació Cellex, Fundació Mir-Puig, Generalitat de Catalunya (CERCA programme), the AXA Chair in Quantum Information Science, EU Quantera project Veriqtas and PASQUANS2, the National Natural Science Foundation of China under Grant No. 12201618, the European Union Horizon’s 2020 research and innovation programme under the Marie Sklodowska Curie grant agreement No 101031549 (QuoMoDys), the NSF grant OAC-1835443 and the ERC Adv. Grant grant agreement No 885682. Research at the Perimeter Institute for Theoretical Physics is supported by the Government of Canada through the Department of Innovation, Science and Economic Development Canada and by the Province of Ontario through the Ministry of Research, Innovation and Science. VM was supported by the FastQI grant funded by the Institut Quantique Occitan, the PHC Proteus grant 46195TA, the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Actions, grant agreement 813211 (POEMA), by the AI Interdisciplinary Institute ANITI funding, through the French “Investing for the Future PIA3” program under the Grant agreement n ANR-19-PI3A-0004 as well as by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. This work was partially performed using HPC resources from CALMIP (Grant 2023-P23035).
@@ Appendix B SDP reductions by exploiting structure
In this appendix, we further discuss how to implement symmetries of the problem in the SDP algorithm to reduce the number of free variables in the implementation. We start by restating the main problem for completeness.
@@
such that: $\quad(\sigma_{i}^{a})^{2}=1,\quad i=1,\ldots,N;a\in\{x,y,z\},$
@@
$\quad\sigma_{i}^{a}\sigma_{j}^{b}=\sigma_{j}^{b}\sigma_{i}^{a},\quad 1\leq i\neq j\leq N;a,b\in\{x,y,z\}.$
@@
$\min\limits_{\{\langle v^{\dagger}w\rangle\}}$ $\quad\langle H\rangle$ (18)
@@
where $\mathbf{M}_{d}$ is the moment matrix indexed by ${\mathcal{B}}_{d}$ with $[{\mathbf{M}}_{d}]_{vw}=\langle v^{\dagger}w\rangle$. The decision variables of (18) are the moments $\langle v^{\dagger}w\rangle,v,w\in{\mathcal{B}}_{d}$. Note that the equality constraints in (17) give rise to the following replacement rules on monomials:
@@ B.2.2 Sign symmetry of the Hamiltonian
Besides the zero entries given in Section B.2.1, these additional sign symmetries of the Hamiltonian yield extra zero entries of the moment matrix: $\langle u\rangle=0$ if $\hbox{\rm{NF}}(u)$ is variant under the transformations (21).
Note that symmetry reduction usually applies to any convex problem with symmetric objective and symmetric feasible set. In this case, the symmetry of the feasible set is not apparent because the replacement rules are not symmetric, e.g., replacing $\sigma_{i}^{x}\sigma_{i}^{y}$ by $\hbox{\bf{i}}\sigma_{i}^{z}$ is not symmetric under any of the substitutions (21). We can show the symmetry of the set of sums of Hermitian squares using (1) their connection with the set of strictly positive Hermitian elements over the quotient ring (detailed below) and (2) the fact that the Hamiltonian is invariant under the substitutions (21).
Indeed, let us denote the non-commutative polynomial ring by ${\mathbb{C}}\langle\{\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z}\}_{i=1}^{N}\rangle$ and the ideal generated by the equality constraints of (17) by $I$. Let $\Omega\coloneqq\{\mathrm{NF}(u)\mid u\text{ is a monomial in }\{\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z}\}_{i=1}^{N}\}$. Then the optimization problem (17) is equivalent to the unconstrained optimization problem: $\min\limits_{\{|\psi\rangle,\sigma_{i}^{a}\}}\langle\psi|H|\psi\rangle$, considered in the quotient ring ${\mathbb{C}}\langle\{\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z}\}_{i=1}^{N}\rangle/I\cong{\mathbb{C}}\langle\Omega\rangle$. Let us denote by $S$ the group of additional sign symmetries given by (21) and by $\Sigma_{S}$ the set of sums of Hermitian squares of ${\mathbb{C}}\langle\Omega\rangle$, that are invariant under $S$ after conversion to normal form, i.e., elements of the form $h=\sum_{j}p_{j}^{\dagger}p_{j}$, $p_{j}\in{\mathbb{C}}\langle\Omega\rangle$, such that $s(\hbox{\rm{NF}}(h))=\hbox{\rm{NF}}(h)$ for any $s\in S$. Then one can show that any strictly positive Hermitian element of ${\mathbb{C}}\langle\Omega\rangle$ that is invariant under $S$ lies in $\Sigma_{S}$ by [10], the proof being very similar to the one of [71, Proposition 3.1]. Here “strict positivity” should be understood as strict positivity over all possible evaluations in Hilbert spaces, as detailed, e.g., in [71, Section 2.2]. By duality one considers optimization over linear functionals nonnegative on $\Sigma_{S}$, which leads to an equivalent formulation of the above unconstrained optimization problem:
@@
To conclude, at each relaxation (18) we can restrict ourselves to optimizing over linear functionals vanishing on variant elements of ${\mathbb{C}}\langle\Omega\rangle$ under the transformations (21). This boils down to setting every entry of the Hermitian moment matrix ${\mathbf{M}}=[\ell(v^{\dagger}w)]_{v,w\in\Omega}$ from (18) to $\langle u\rangle=0$ if $\hbox{\rm{NF}}(u)$ is variant under the transformations (21).